A strip of uniform width is to be cut off of both sides and both ends of a sheet of paper that is 20 inches by 6 inches, in order to reduce the size of the paper to an area of 32 square inches. Find the width of the strip.

Answer by cleomenius(959)

20 * 6 = 120
(20 - x)(6 - x) = 32
120 - 26x +x^2 = 32
x^2 - 26x + 120 = 32
x^2 -26x + 88 = 0
(x - 4)(x + 22)
x = 4
16 * 2 = 32 inches.